# Physical Processes Behind the Co-Evolution of Halos, Galaxies and Supermassive Black Holes in the IllustrisTNG Simulation

### v2

@@ 5.2 Inferring the paths of SMBH growth from observations
Figure 12: Distribution and movement of galaxies in the $M_{\rm BH}$-$M_{*}$ plane. 2-D histogram is shown for all central galaxies in TNG with $M_{*}\geqslant 10^{8.5}\,{\rm M_{\odot}}$ at $z=0$. Colors are encoded and contours are drawn to show their host halo masses. Red arrows are path segments of these galaxies from $z=0.3$ to $z=0$. Tilted dashed line shows the $M_{\rm BH}$-$M_{\rm*,bulge}$ relation found in observation by Kormendy & Ho (2013). See §5.2 for a detailed discussion.
@@
Fig. 12 shows the distribution of SMBHs in the $M_{\rm BH}$-$M_{*}$ plane at $z=0$, with colors and contours encoded according to their host halo masses. For each population of SMBHs with a given $M_{\rm BH}$ and $M_{*}$, an arrow is drawn to indicate the movement of its location in this plane since $z=0.3$. To cover larger ranges of $M_{\rm BH}$ and $M_{*}$ for this narrow redshift interval, we enlarge the sample size by including all branches with $M_{*,z=0}\geqslant 10^{8.5}\,{\rm M_{\odot}}$. Clear trends can be seen in different regions, each containing SMBHs in a given phase discussed in §4. The lower left corner of the plane is populated with SMBHs in Phase-1, where $M_{\rm BH}$ is close to the seed mass of $\sim 10^{6}\,{\rm M_{\odot}}$. The arrows in the region of $M_{*}\sim 10^{9.5}\,{\rm M_{\odot}}$ and $M_{\rm BH}\lesssim 10^{7}\,{\rm M_{\odot}}$ aligns with the vertical axis, which marks the transition from the star-formation dominated growth to the SMBH-accretion dominated growth (Phase-2). Once SMBHs approach the local scaling relation of $M_{\rm BH}$-$M_{*}$ from either the upper or lower side, the flow moves along this relation by the self-regulation dominated growth (Phase-3). The prevalence of massive central SMBHs of $M_{\rm BH}\sim 10^{8.2}\,{\rm M_{\odot}}$ (Terrazas et al., 2020, see their figure 5) drives the quenching of galaxies and freezes the evolution of the SMBHs in the $M_{\rm BH}$-$M_{*}$ plane during Phase-4. For the most massive SMBHs, mergers tend to drive a “dry” evolution towards higher masses, although the sample size is not enough at the high-mass end to produce stable flow lines.
@@
Drawing the flow lines in the $M_{\rm BH}$-$M_{\rm*}$ plane from observations is more challenging, as it requires the measurements of four quantities $M_{\rm BH}$, $M_{*}$, BHAR and SFR. For a galaxy with inactive SMBH, BHAR is too low to be detected, while for a galaxy with active SMBH, a decomposition of its image and spectrum is needed to separate the contributions from the AGN and stars. The available data thus only provide an incomplete view of the SMBH populations, together with measurement uncertainties that probably mix flow lines. Thus, caution should be taken when interpreting the results in order to avoid biased conclusions. A recent work by Zhuang & Ho (2023, see their figure 4) indeed achieved this goal by stacking a large sample of AGNs at $z\leqslant 0.35$, with physical properties derived by a simultaneous decomposition of image and spectrum. The set of flow lines they obtained converges to the local scaling relation, thus supporting the results produced by the subgrid model of TNG for the SMBH growth. A noticeable difference between their results and those of TNG is the presence of over-massive black holes lying above the local scaling relation. This difference is either due to uncertainties in their measurements, or due to the strong feedback in TNG that forbids the growth of SMBHs beyond the local scaling relation. Future observations and more precise measurements are needed to pin down the exact reason.
@@
An observationally more accessible way to infer the evolutionary paths is to only use measurements of the mass of different components. One possibility is to use the size of the scatter in the $M_{\rm BH}$-$M_{\rm*}$ plane to distinguish different phases. This has been suggested by McAlpine et al. (2018), who found that the EAGLE simulation produces a redshift-dependent critical halo mass for the SMBHs to transit from Phase-1 to Phase-2. As Phase-2 has rapid $M_{\rm BH}$ growth while Phase-1 and Phase-3 do not, the scatter in $M_{\rm BH}$ is expected to be large during Phase-2 relative to its two sides where the slopes of the $M_{\rm BH}$-$M_{*}$ and $M_{\rm BH}$-$M_{\rm h}$ relations are shallower. Another possibility is to use multivariate statistical methods, such as the regression of $M_{\rm BH}$ on the joint of $M_{*}$ and $M_{\rm h}$. This can be seen from the contours in Fig. 12: for $M_{\rm h}<10^{12.3}\,{\rm M_{\odot}}$ (or $M_{*}<10^{10.5}\,{\rm M_{\odot}}$), corresponding to $M_{\rm BH}<10^{8}\,{\rm M_{\odot}}$, the contours are almost vertical, indicating a strong correlation between $M_{\rm BH}$ and $M_{\rm*}$. At higher masses, the contours become horizontal, suggesting that the growth of the SMBH is related more to the halo than to the galaxy. Such a change is implied by the transition from Phase-3 (§4.3), where the growth of SMBH is regulated by the thermal-mode feedback that balances the gravity of gas within the galaxy, to Phase-4 (§4.4), where the growth of SMBH is reduced by the kinetic-mode feedback that balances the cooling of the whole central subhalo. Thus, the scale on which the feedback energy is coupled to the surrounding gas is imprinted in the correlation between mass variables. Indeed, Zhang et al. (2024) found an observational trend similar to the contours shown in Fig. 12, using $M_{\rm h}$ measured from galaxy-galaxy weak lensing and $M_{\rm BH}$ converted from stellar velocity dispersion. The transition mass scales they obtained, $M_{\rm h}\sim 10^{12.0}\,{\rm M_{\odot}}$ and $M_{\rm BH}\sim 10^{7.4}\,{\rm M_{\odot}}$, are consistent with our findings from TNG.
@@ 6.1 Main results
[…]
@@
    Although the growth histories of individual galaxies in the $M_{\rm BH}$-$M_{*}$ plane are diverse, they show regularities represented by four distinct phases. We fit the $M_{\rm BH}$-$M_{*}$ history of each galaxy in our sample with a function that allows us to separate the four phases and locate the transition points between them. We found that the first two phases are short, about $1$-$2\,{\rm Gyr}$ on average, while the last two phases are significantly longer. Each phase is found to be associated with a process that dominates the growth of the SMBH and its host galaxy, and physical conditions can be identified for the transition from one phase to the next (§3).
@@
    The power index ($\gamma$) of the $M_{\rm BH}$-$M_{*}$ relation varies as the dominating processes for the growth of SMBHs and galaxies change between different phases. The value of $\gamma$ is about $1.5$ at $t_{23}$ due to the condition for the thermal-mode feedback to overcome the potential energy of the gas; it becomes smaller than $1$ at $t_{\rm q}$ due to the specific choice of TNG for triggering the kinetic-mode feedback; and it is driven back to $\approx 1$ at $z=0$ for high-mass galaxies by mergers. We highlight potential observational tests for the predictions of the simulation. These include the use of scaling relations that bracket the phases and transitions, and the use of the mass growth rates to reconstruct the evolution paths of SMBHs and galaxies (§5).
@@ 6.2 Implications
We note that our results and conclusions are based on the numerical and subgrid models of IllustrisTNG. How the predicted observables respond to the change of model assumptions is of great interest and is one of the main targets in using hydrodynamical simulations to understand galaxy formation. Such a question has been explored by Habouzit et al. (2021) using a set of different simulations: Illustris, IllustrisTNG, Horizon-AGN, EAGLE and SIMBA. All these simulations produce a $M_{\rm BH}$-$M_{*}$ relation similar to that observed for high-mass ($M_{*}\gtrsim 10^{10.5}\,{\rm M_{\odot}}$) galaxies at $z\sim 0$ (e.g. Kormendy & Ho, 2013). Such an agreement is not significant, however, as the observational relation was used to calibrate the model parameters in these simulations. More interesting and significant is that the paths of individual galaxies in the $M_{\rm BH}$-$M_{*}$ plane before they reach the local scaling relation are different both among different simulations and at different redshifts within a given simulation.
@@
In EAGLE, the $M_{\rm BH}$-$M_{*}$ relation at both $z=3$ and $z=0$ is found to be non-linear (in logarithmic scale) and to drop at $M_{*}\lesssim 10^{10.5}\,{\rm M_{\odot}}$, reflecting the strong supernova feedback in suppressing the growth of SMBHs (corresponding to our Phase-1) in low-mass galaxies at both redshifts. As discussed in §4.1 and based on the results of McAlpine et al. (2017), the transition from Phase-1 to Phase-2 in EAGLE occurs when the virial temperature of the halo reaches a time-independent critical value. The corresponding critical halo mass changes moderately from $M_{\rm h}\approx 10^{11.2}\,{\rm M_{\odot}}$ at $z=6$ to $10^{12.4}\,{\rm M_{\odot}}$ at $z=0$. At a given redshift, the observed $M_{\rm BH}$-$M_{*}$ relation at the low-mass end is thus dominated by the insufficient growth of SMBHs in Phase-1, which causes the relation to bend downwards with respect to the local scaling relation.
@@
The situation in TNG is different in that the non-linearity of the $M_{\rm BH}$-$M_{*}$ relation at the low-mass end appears at $z=3$ but disappears at $z=0$. Habouzit et al. (2021) attributed this to the “inefficient” SN feedback at low $z$ in TNG. We can follow the argument used in deriving $\tau_{1}$ in §4.1 to obtain some quantitative understanding, by explicitly incorporating the definition of the transition into the Bondi accretion. Equaling the sBHAR given by the Bondi formula (Eq. 3) to the sSFR, we obtain the SMBH mass at the transition $t_{12}$ as $M_{\rm BH}(t_{12})={\rm sSFR}(t_{12})/A_{\rm s}(t_{12})$. Because $A_{\rm s}$ can be approximated as a constant during Phase-1 (see Fig. 3) and the sSFR of star-forming galaxies decreases by $\gtrsim 1\,{\rm dex}$ from $z=3$ to $z=0$ in TNG (e.g. Chen et al., 2021, see their figure 8), $M_{\rm BH}(t_{12})$ is expected to decrease by $\gtrsim 1\,{\rm dex}$ in this redshift range. As seen from Fig. 2, this decrease is sufficient to push $M_{\rm BH}(t_{12})$ downwards to a position below the seed mass of SMBHs. Thus, a newly seeded SMBH at low $z$ skips the entire Phase-1, which makes the $M_{\rm BH}$-$M_{*}$ relation linear towards the low-mass end. Indeed, for galaxies at $z\leqslant 0.3$ shown in Fig. 12, the flow lines at $M_{\rm BH}\approx 10^{6}\,{\rm M_{\odot}}$ go upwards without a significant horizontal movement, as expected from our argument. Our result also implies that, if a lower seed mass is adopted while keeping other recipes unchanged in TNG, or if the seeding is delayed until a higher halo/stellar mass is reached, the non-linear $M_{\rm BH}$-$M_{*}$ relation would appear at lower $z$. The latter condition is fulfilled in SIMBA, and a non-linear $M_{\rm BH}$-$M_{*}$ relation towards the low-mass end is indeed found at $z=0$ in this simulation (Habouzit et al., 2021, see their figure 11).
@@
In more extreme cases where SN feedback is weak even at high $z$ or totally absent in the simulation, a newly seeded SMBH would immediately start its rapid growth and Phase-1 is expected to disappear completely. The weak SN feedback adopted by Horizon-AGN is such a case, and indeed the predicted $M_{\rm BH}$-$M_{*}$ relation at $z=3$ is roughly linear (Habouzit et al., 2021, see their figure 11). Numerical experiments of disabling SN feedback by Bower et al. (2017, see their figure 7) also produce such a linear relation.
@@
The duration of Phase-2, $\tau_{2}$, as shown in Fig. 6 and estimated by Eq. (4.1), is about $1\,{\rm Gyr}$. However, this time scale depends on the gas environment around a SMBH, and thus depends on the density threshold of star-forming gas assumed by the simulation, the subgrid models evolving the cold gas, and the smoothing kernel defining the SMBH neighborhood. The threshold density is $n_{\rm gas}\approx 0.1\,{\rm cm}^{-3}$ in both EAGLE and TNG, far below the density that a real galaxy can reach, as inferred both analytically and observationally (see, e.g. §2 of Chen et al. 2024b and §6 of Dekel et al. 2023). Lifting this threshold is computationally expensive and thus not feasible for simulations on cosmological scales. Using a set of idealized simulations for individual gas clouds embedded with randomly seeded low-mass BHs, Shi et al. (2023) found that the small-scale structure of a gas cloud can be complex. The runaway collision of a BH with gas sub-cloud produces a boost of $M_{\rm BH}$, much resembling the rapid growth of Phase-2. The collision can repeat multiple times, leading to a ladder-like growth history of $M_{\rm BH}$ (see their figures 1 and 2). Each collision starts and ends within a duration $\ll 1\,{\rm Myr}$, more than three orders of magnitude shorter than our $\tau_{2}$, and even shorter than the lifetime of massive stars so that SNs are not able to be triggered. On the other hand, Hopkins et al. (2024b) conducted a “hyper-refinement” zoom-in simulation for a galaxy within the cosmological context. Asymmetric gas structures around the SMBH are found and sustained super-Eddington accretion flow coexists with outflow. Star-formation within the influence radius of the SMBH is suppressed due to the pressure from turbulence and magnetic field, making star formation and stellar feedback ineffective. The high density and complex sub-structure of gas, the fast and repeated boost of $M_{\rm BH}$ by super- or hyper-Eddington accretion within a time scale less than $\,{\rm Myr}$, and the inefficient stellar feedback, all suggest rich physics of Phase-2 that needs to be explored further.
@@
The ending of Phase-2 at $t_{23}$ by the AGN feedback to unbind the gas (Eq. 22) depends on the (radiative and coupling) efficiency of the feedback energy ($\eta_{\rm thermal}$) and subsequent cooling of the heated gas ($\eta_{\rm eff}$). A simulation with different $\eta_{\rm thermal}$ and/or $\eta_{\rm eff}$ can thus produce a $M_{\rm BH}$-$M_{*}$ relation at $t_{23}$ (Fig. 11) with different normalizations. Since the self-regulation of SMBHs during Phase-3 is also determined by these two factors, the normalization of the $M_{\rm BH}$-$M_{*}$ relation during Phase-3 and at $t_{\rm q}$ will also be changed. The difference in the normalization of the scaling relation shown in figure 4 of Habouzit et al. (2021) for different simulations can thus be understood along this line. On the other hand, the power index of the scaling relation at a specific event such as $t_{23}$ or $t_{\rm q}$, is distorted by the conditions for the event to be triggered. We showed in §5.1 that the kinetic-mode AGN feedback modeled in TNG is preferentially triggered in high-mass halos, which produces a shallow power index of the scaling relation at $t_{\rm q}$. Such a transition in SIMBA does not have a clear mass preference, and thus produces a more widely spread range of $M_{*}$ and a non-constant power index.
The final phase, Phase-4, is dominated by the ex-situ, gas-poor growth at the high-mass end. As the frequency of mergers is determined by the hierarchical structure formation in the $\Lambda$CDM cosmology, it is not sensitive to baryonic physics and is thus expected to be similar across simulations. Indeed, all the simulations shown in figure 4 of Habouzit et al. (2021) produce a tip at the high-mass end of the $M_{\rm BH}$-$M_{*}$ relation, with a power index of $\approx 1$. The normalization of this tip, however, inherits the influences of baryonic physics at $t_{\rm q}$, and thus varies with simulations, as seen from the lower normalization of EAGLE. The scaling relation in Phase-4 for low-mass galaxies that have not experienced significant mergers is frozen after $t_{\rm q}$, and thus also depends on the baryonic physics.
@@
Data directly related to this publication and its figures are available on request from the corresponding author. The IllustrisTNG simulation themselves, are publicly available and accessible at www.tng-project.org/data.
This work is supported by the National Natural Science Foundation of China (NSFC, Nos. 12192224, and 11890693), CAS Project for Young Scientists in Basic Research (No. YSBR-062), the Fundamental Research Funds for the Central Universities, and the China Postdoctoral Science Foundation (No. 2022TQ0329). We acknowledge the science research grants from the China Manned Space Project with No. CMS-CSST-2021-A03. The authors gratefully acknowledge the support of Cyrus Chun Ying Tang Foundations. HL thanks Kai Wang, Hui Hong, Xiong Luo, Yuan Wang, Enci Wang, Wentao Luo and Yu Rong for their valuable insights and discussions. The work is also supported by the Supercomputer Center of University of Science and Technology of China. The authors would like to express their gratitude to Max Planck Computing and Data Facility (MPCDF) for providing the necessary computational and data storage resources that have significantly contributed to the research results presented in this paper.

### v3

@@ 5.2 Inferring the paths of SMBH growth from observations
Figure 12: Distribution and movement of galaxies in the $_{\rm BH}-MM_{*}$ plane. 2-D histogram is shown for all central galaxies in TNG with $M_{*}\geqslant 10^{8.5}\,{\rm M_{\odot}}$ at $z=0$. Colors are encoded and contours are drawn to show their host halo masses. Red arrows are path segments of these galaxies from $z=0.3$ to $z=0$. Tilted dashed line shows the $M_{\rm BH}$-$M_{\rm*,bulge}$ relation found in observation by Kormendy & Ho (2013). See §5.2 for a detailed discussion.
@@
Fig. 12 shows the distribution of SMBHs in the $_{\rm BH}-MM_{*}$ plane at $z=0$, with colors and contours encoded according to their host halo masses. For each population of SMBHs with a given $M_{\rm BH}$ and $M_{*}$, an arrow is drawn to indicate the movement of its location in this plane since $z=0.3$. To cover larger ranges of $M_{\rm BH}$ and $M_{*}$ for this narrow redshift interval, we enlarge the sample size by including all branches with $M_{*,z=0}\geqslant 10^{8.5}\,{\rm M_{\odot}}$. Clear trends can be seen in different regions, each containing SMBHs in a given phase discussed in §4. The lower left corner of the plane is populated with SMBHs in Phase-1, where $M_{\rm BH}$ is close to the seed mass of $\sim 10^{6}\,{\rm M_{\odot}}$. The arrows in the region of $M_{*}\sim 10^{9.5}\,{\rm M_{\odot}}$ and $M_{\rm BH}\lesssim 10^{7}\,{\rm M_{\odot}}$ align with the vertical axis, which marks the transition from the star-formation dominated growth to the SMBH-accretion dominated growth (Phase-2). Once SMBHs approach the local scaling relation of $_{\rm BH}-MM_{*}$ from either the upper or lower side, the flow moves along this relation by the self-regulation dominated growth (Phase-3). The prevalence of massive central SMBHs of $M_{\rm BH}\sim 10^{8.2}\,{\rm M_{\odot}}$ (Terrazas et al., 2020, see their figure 5) drives the quenching of galaxies and freezes the evolution of the SMBHs in the $_{\rm BH}-MM_{*}$ plane during Phase-4. For the most massive SMBHs, mergers tend to drive a “dry” evolution towards higher masses, although the sample size is not enough at the high-mass end to produce stable flow lines.
@@
Drawing the flow lines in the $_{\rm BH}-MM_{\rm*}$ plane from observations is more challenging, as it requires the measurements of four quantities $_{\rm BH}, MM_{*}$, BHAR and SFR. For a galaxy with inactive SMBH, BHAR is too low to be detected, while for a galaxy with active SMBH, a decomposition of its image and spectrum is needed to separate the contributions from the AGN and stars. The available data thus only provide an incomplete view of the SMBH populations, together with measurement uncertainties that probably mix flow lines. Thus, caution should be taken when interpreting the results in order to avoid biased conclusions. A recent work by Zhuang & Ho (2023, see their figure 4) indeed achieved this goal by stacking a large sample of AGNs at $z\leqslant 0.35$, with physical properties derived by a simultaneous decomposition of image and spectrum. The set of flow lines they obtained converges to the local scaling relation, thus supporting the results produced by the subgrid model of TNG for the SMBH growth. A noticeable difference between their results and those of TNG is the presence of over-massive black holes lying above the local scaling relation. This difference is either due to uncertainties in their measurements, or due to the strong feedback in TNG that forbids the growth of SMBHs beyond the local scaling relation. Future observations and more precise measurements are needed to pin down the exact reason.
@@
An observationally more accessible way to infer the evolutionary paths is to only use measurements of the mass of different components. One possibility is to use the size of the scatter in the $_{\rm BH}-MM_{\rm*}$ plane to distinguish different phases. This has been suggested by McAlpine et al. (2018), who found that the EAGLE simulation produces a redshift-dependent critical halo mass for the SMBHs to transit from Phase-1 to Phase-2. As Phase-2 has rapid $M_{\rm BH}$ growth while Phase-1 and Phase-3 do not, the scatter in $M_{\rm BH}$ is expected to be large during Phase-2 relative to its two sides where the slopes of the $_{\rm BH}-MM_{*}$ and $M_{\rm BH}$-$M_{\rm h}$ relations are shallower. Another possibility is to use multivariate statistical methods, such as the regression of $M_{\rm BH}$ on the joint of $M_{*}$ and $M_{\rm h}$. This can be seen from the contours in Fig. 12: for $M_{\rm h}<10^{12.3}\,{\rm M_{\odot}}$ (or $M_{*}<10^{10.5}\,{\rm M_{\odot}}$), corresponding to $M_{\rm BH}<10^{8}\,{\rm M_{\odot}}$, the contours are almost vertical, indicating that $M_{\rm BH}$ does not depend on $M_{\rm h}$ once $M_{*}$ is fixed. At higher masses, the contours become horizontal, suggesting that the growth of the SMBH is related more to the halo than to the galaxy. Such a change is implied by the transition from Phase-3 (§4.3), where the growth of SMBH is regulated by the thermal-mode feedback that balances the gravity of gas within the galaxy, to Phase-4 (§4.4), where the growth of SMBH is reduced by the kinetic-mode feedback that balances the cooling of the whole central subhalo. Thus, the scale on which the feedback energy is coupled to the surrounding gas is imprinted in the correlation between mass variables. Indeed, Zhang et al. (2024) found an observational trend similar to the contours shown in Fig. 12, using $M_{\rm h}$ measured from galaxy-galaxy weak lensing and $M_{\rm BH}$ converted from stellar velocity dispersion. The transition mass scales they obtained, $M_{\rm h}\sim 10^{12.0}\,{\rm M_{\odot}}$ and $M_{\rm BH}\sim 10^{7.4}\,{\rm M_{\odot}}$, are consistent with our findings from TNG.
@@ 6.1 Main results
In this paper, we used the IllustrisTNG simulation to quantify the co-evolution of SMBHs, their host galaxies, and halos. We focus on galaxies that are quenched at z=0z=02) and use the main branches of the merger trees rooted in these galaxies to trace their evolution. We summarize our results and conclusions in the following.
@@
    Although the growth histories of individual galaxies in the $_{\rm BH}-MM_{*}$ plane are diverse, they show regularities represented by four distinct phases. We fit the $_{\rm BH}-MM_{*}$ history of each galaxy in our sample with a function that allows us to separate the four phases and locate the transition points between them. We found that the first two phases are short, about $1$-$2\,{\rm Gyr}$ on average, while the last two phases are significantly longer. Each phase is found to be associated with a process that dominates the growth of the SMBH and its host galaxy, and physical conditions can be identified for the transition from one phase to the next (§3).
@@
    The power index ($\gamma$) of the $_{\rm BH}-MM_{*}$ relation varies as the dominating processes for the growth of SMBHs and galaxies change between different phases. The value of $\gamma$ is about $1.5$ at $t_{23}$ due to the condition for the thermal-mode feedback to overcome the potential energy of the gas; it becomes smaller than $1$ at $t_{\rm q}$ due to the specific choice of TNG for triggering the kinetic-mode feedback; and it is driven back to $\approx 1$ at $z=0$ for high-mass galaxies by mergers. We highlight potential observational tests for the predictions of the simulation. These include the use of scaling relations that bracket the phases and transitions, and the use of the mass growth rates to reconstruct the evolution paths of SMBHs and galaxies (§5).
@@ 6.2 Implications
We note that our results and conclusions are based on the numerical and subgrid models of IllustrisTNG. How the predicted observables respond to the change of model assumptions is of great interest and is one of the main targets in using hydrodynamical simulations to understand galaxy formation. Such a question has been explored by Habouzit et al. (2021) using a set of different simulations: Illustris, IllustrisTNG, Horizon-AGN, EAGLE and SIMBA. All these simulations produce a $_{\rm BH}-MM_{*}$ relation similar to that observed for high-mass ($M_{*}\gtrsim 10^{10.5}\,{\rm M_{\odot}}$) galaxies at $z\sim 0$ (e.g. Kormendy & Ho, 2013). Such an agreement is not significant, however, as the observational relation was used to calibrate the model parameters in these simulations. More interesting and significant is that the paths of individual galaxies in the $_{\rm BH}-MM_{*}$ plane before they reach the local scaling relation are different both among different simulations and at different redshifts within a given simulation.
@@
In EAGLE, the $_{\rm BH}-MM_{*}$ relation at both $z=3$ and $z=0$ is found to be non-linear (in logarithmic scale) and to drop at $M_{*}\lesssim 10^{10.5}\,{\rm M_{\odot}}$, reflecting the strong supernova feedback in suppressing the growth of SMBHs (corresponding to our Phase-1) in low-mass galaxies at both redshifts. As discussed in §4.1 and based on the results of McAlpine et al. (2017), the transition from Phase-1 to Phase-2 in EAGLE occurs when the virial temperature of the halo reaches a time-independent critical value. The corresponding critical halo mass changes moderately from $M_{\rm h}\approx 10^{11.2}\,{\rm M_{\odot}}$ at $z=6$ to $10^{12.4}\,{\rm M_{\odot}}$ at $z=0$. At a given redshift, the observed $_{\rm BH}-MM_{*}$ relation at the low-mass end is thus dominated by the insufficient growth of SMBHs in Phase-1, which causes the relation to bend downwards with respect to the local scaling relation.
@@
The situation in TNG is different in that the non-linearity of the $_{\rm BH}-MM_{*}$ relation at the low-mass end appears at $z=3$ but disappears at $z=0$. Habouzit et al. (2021) attributed this to the “inefficient” SN feedback at low $z$ in TNG. We can follow the argument used in deriving $\tau_{1}$ in §4.1 to obtain some quantitative understanding, by explicitly incorporating the definition of the transition into the Bondi accretion. Equaling the sBHAR given by the Bondi formula (Eq. 3) to the sSFR, we obtain the SMBH mass at the transition $t_{12}$ as $M_{\rm BH}(t_{12})={\rm sSFR}(t_{12})/A_{\rm s}(t_{12})$. Because $A_{\rm s}$ can be approximated as a constant during Phase-1 (see Fig. 3) and the sSFR of star-forming galaxies decreases by $\gtrsim 1\,{\rm dex}$ from $z=3$ to $z=0$ in TNG (e.g. Chen et al., 2021, see their figure 8), $M_{\rm BH}(t_{12})$ is expected to decrease by $\gtrsim 1\,{\rm dex}$ in this redshift range. As seen from Fig. 2, this decrease is sufficient to push $M_{\rm BH}(t_{12})$ downwards to a position below the seed mass of SMBHs. Thus, a newly seeded SMBH at low $z$ skips the entire Phase-1, which makes the $_{\rm BH}-MM_{*}$ relation linear towards the low-mass end. Indeed, for galaxies at $z\leqslant 0.3$ shown in Fig. 12, the flow lines at $M_{\rm BH}\approx 10^{6}\,{\rm M_{\odot}}$ go upwards without a significant horizontal movement, as expected from our argument. Our result also implies that, if a lower seed mass is adopted while keeping other recipes unchanged in TNG, or if the seeding is delayed until a higher halo/stellar mass is reached, the non-linear $_{\rm BH}-MM_{*}$ relation would appear at lower $z$. The latter condition is fulfilled in SIMBA, and a non-linear $_{\rm BH}-MM_{*}$ relation towards the low-mass end is indeed found at $z=0$ in this simulation (Habouzit et al., 2021, see their figure 11).
@@
In more extreme cases where SN feedback is weak even at high $z$ or totally absent in the simulation, a newly seeded SMBH would immediately start its rapid growth and Phase-1 is expected to disappear completely. The weak SN feedback adopted by Horizon-AGN is such a case, and indeed the predicted $_{\rm BH}-MM_{*}$ relation at $z=3$ is roughly linear (Habouzit et al., 2021, see their figure 11). Numerical experiments of disabling SN feedback by Bower et al. (2017, see their figure 7) also produce such a linear relation.
@@
The duration of Phase-2, $\tau_{2}$, as shown in Fig. 6 and estimated by Eq. (4.1), is about $1\,{\rm Gyr}$. However, this timescale depends on the gas environment around a SMBH, and thus depends on the density threshold of star-forming gas assumed by the simulation, the subgrid models evolving the cold gas, and the smoothing kernel defining the SMBH neighborhood. The threshold density is $n_{\rm gas}\approx 0.1\,{\rm cm}^{-3}$ in both EAGLE and TNG, far below the density that a real galaxy can reach, as inferred both analytically and observationally (see, e.g. §2 of Chen et al. 2024b and §6 of Dekel et al. 2023). Lifting this threshold is computationally expensive and thus not feasible for simulations on cosmological scales. Using a set of idealized simulations for individual gas clouds embedded with randomly seeded low-mass BHs, Shi et al. (2023) found that the small-scale structure of a gas cloud can be complex. The runaway collision of a BH with a gas sub-cloud produces a boost of $M_{\rm BH}$, much resembling the rapid growth of Phase-2. The collision can repeat multiple times, leading to a ladder-like growth history of $M_{\rm BH}$ (see their figures 1 and 2). Each collision starts and ends within a duration $\ll 1\,{\rm Myr}$, more than three orders of magnitude shorter than our $\tau_{2}$, and even shorter than the lifetime of massive stars so that SNs are not able to be triggered. On the other hand, Hopkins et al. (2024) conducted a “hyper-refinement” zoom-in simulation for a galaxy within the cosmological context. Asymmetric gas structures around the SMBH are found and sustained super-Eddington accretion flow coexists with outflow. Star-formation within the influence radius of the SMBH is suppressed due to the pressure from turbulence and magnetic field, making star formation and stellar feedback ineffective. The high density and complex sub-structure of gas, the fast and repeated boost of $M_{\rm BH}$ by super- or hyper-Eddington accretion within a timescale less than $\rm Myr$}, and the inefficient stellar feedback, all suggest rich physics of Phase-2 that needs to be explored further.
@@
The ending of Phase-2 at $t_{23}$ by the AGN feedback to unbind the gas (Eq. 22) depends on the (radiative and coupling) efficiency of the feedback energy ($\eta_{\rm thermal}$) and subsequent cooling of the heated gas ($\eta_{\rm eff}$). A simulation with different $\eta_{\rm thermal}$ and/or $\eta_{\rm eff}$ can thus produce a $_{\rm BH}-MM_{*}$ relation at $t_{23}$ (Fig. 11) with different normalizations. Since the self-regulation of SMBHs during Phase-3 is also determined by these two factors, the normalization of the $_{\rm BH}-MM_{*}$ relation during Phase-3 and at $t_{\rm q}$ will also be changed. The difference in the normalization of the scaling relation shown in figure 4 of Habouzit et al. (2021) for different simulations can thus be understood along this line. On the other hand, the power index of the scaling relation at a specific event such as $t_{23}$ or $t_{\rm q}$, is distorted by the conditions for the event to be triggered. We showed in §5.1 that the kinetic-mode AGN feedback modeled in TNG is preferentially triggered in high-mass halos, which produces a shallow power index of the scaling relation at $t_{\rm q}$. Such a transition in SIMBA does not have a clear mass preference, and thus produces a more widely spread range of $M_{*}$ and a non-constant power index.
[…]
@@
Data directly related to this publication and its figures are available on request from the corresponding author. The IllustrisTNG simulations themselves, are publicly available and accessible at www.tng-project.org/data.
We thank the referee for a useful report that significantly improved the paper. This work is supported by the National Natural Science Foundation of China (NSFC, Nos. 12192224, and 11890693), CAS Project for Young Scientists in Basic Research (No. YSBR-062), the Fundamental Research Funds for the Central Universities, and the China Postdoctoral Science Foundation (No. 2022TQ0329). We acknowledge the science research grants from the China Manned Space Project with No. CMS-CSST-2021-A03. The authors gratefully acknowledge the support of Cyrus Chun Ying Tang Foundations. HL thanks Kai Wang, Hui Hong, Xiong Luo, Yuan Wang, Enci Wang, Wentao Luo and Yu Rong for their valuable insights and discussions. The work is also supported by the Supercomputer Center of University of Science and Technology of China. The authors would like to express their gratitude to Max Planck Computing and Data Facility (MPCDF) for providing the necessary computational and data storage resources that have significantly contributed to the research results presented in this paper.